OpenStudy (anonymous):

Please explain to me how the adding of two negative numbers equals a positive number. When they were shoving number lines and set subset theory at me, this issue came up and I never got a good answer.

OpenStudy (anonymous):

eg. -2-2=-4 when two negative sign comes we want to add and put the negative sign because when the sign are same we wanted to put that sign itself eg. +2+2=+2

OpenStudy (anonymous):

if u are adding positive and negative eg -5+3=-2 when opposite sign come we want to subract and put the greater numbers sign -5 is greater here so that we put negative sign

OpenStudy (anonymous):

now cleared

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

this is easy if we understand that

OpenStudy (anonymous):

ya true